# Massless Quasiparticles in Bogoliubov-de Gennes Systems

### v1

@@ F.3 Consequences of $\mathcal{C}$ symmetry
Firstly, we consider a special case in which $\mathcal{C}$ preserves the supporting space of every minimal nonzero modes $\mathcal{L}_{\varepsilon_{k}}\oplus\mathcal{L}_{-\varepsilon_{k}}$, namely we first assume that $\mathcal{C}$ does not cause additional degeneracy for the energy eigenstates with $\varepsilon_{k}\not=0$. Such charge-conjugation symmetries all have the following property: for any given minimal nonzero mode, there exists an extension from $F_{f}(k)$ to $F^{c}_{f}(k)$. Recall that a set of minimal nonzero modes either
(I) form an irrep of $F_{f}(k)$ (we denote it as $D_{a}(F_{f}(k))$ whereas the restricted rep $D_{a}(G_{f}(k))=m_{a}(G_{f}(k))\oplus\tilde{m}_{a}(G_{f}(k))$ is reducible), or
@@
(II) form a direct sum of two irreps of $F_{f}(k)$ (we denote it as $D_{a}(F_{f}(k))=\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$).
@@
Then the presence of $\mathcal{C}$ can affect the set of minimal nonzero modes in different ways. Here we discuss them separately.
(I) For the irreducible nonzero modes $D_{a}(F_{f}(k))$.
@@
1) If the irrep $D_{a}(G_{f}^{c}(k))$ is an induction from $m_{a}(G_{f}(k))$ (or from $\tilde{m}_{a}(G_{f}(k))$), namely, there does not exist any extension from $m_{a}(G_{f}(k))$ or $\tilde{m}_{a}(G_{f}(k))$ to $G_{f}^{c}(k)$, then $D_{a}(F^{c}_{f}(k))$ corresponds a set of irreducible zero modes of $F_{f}^{c}(k)$.
@@
2) If there exists an extension from $m_{a}(G_{f}(k))$ (or $\tilde{m}_{a}(G_{f}(k)))$ to $m_{a}(G^{c}_{f}(k))$ (or $\tilde{m}_{a}(G^{c}_{f}(k))$), then $D_{a}(F^{c}_{f}(k))$ remains to be an irreducible nonzero mode of $F_{f}^{c}(k)$.
(II) For the reducible minimal nonzero modes $D_{a}(F_{f}(k))=\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$.
@@
1) If there exist an extension from $\mathcal{D}_{\pm}(F_{f}(k))$ to $\mathcal{D}_{\pm}(F_{f}^{c}(k))$ (hence there must exist an extension from $m_{a}(G_{f}(k))$ to $m_{a}(G_{f}^{c}(k))$), then $\mathcal{D}_{\pm}(F^{c}_{f}(k))$ will form two sets of irreducible zero modes of $F_{f}^{c}(k)$. If $\mathcal{D}_{\pm}(F^{c}_{f}(k))$ satisfy the conditions listed in Tab.1, namely if they can couple to each other, then $\mathcal{D}_{+}(F^{c}_{f}(k))\oplus\mathcal{D}_{-}(F^{c}_{f}(k))$ remains to be a set of reducible minimal nonzero modes; on the other hand, if they cannot couple to each other, (for example, $\mathcal{C}$ provides different quantum numbers in $\mathcal{D}_{+}(F^{c}_{f}(k))$ and $\mathcal{D}_{-}(F^{c}_{f}(k))$ respectively to prevent them from coupling to each other), then $\mathcal{D}_{+}(F^{c}_{f}(k))\oplus\mathcal{D}_{-}(F^{c}_{f}(k))$ will form zero modes.
@@
2) If $\mathcal{C}$ couples $\mathcal{D}_{+}(F_{f}(k))$ with $\mathcal{D}_{-}(F_{f}(k))$ such that the direct sum space $\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$ forms an irrep $D(F_{f}^{c}(k))$ for $F_{f}^{c}(k)$ (namely $D(F_{f}^{c}(k))$ is an induction form $\mathcal{D}_{+}(F_{f}(k))$ to $F_{f}^{c}(k)$) and if there is no extension from $m_{a}(G_{f}(k))$ to $G_{f}^{c}(k)$, then $D(F_{f}^{c}(k))$ forms irreducible zero modes of $F_{f}^{c}(k)$.
@@
3) If $\mathcal{C}$ couples $\mathcal{D}_{+}(F_{f}(k))$ with $\mathcal{D}_{-}(F_{f}(k))$ such that the direct sum space $\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$ forms an irrep $D(F_{f}^{c}(k))$ for $F_{f}^{c}(k)$ and if there exists extension from $m_{a}(G_{f}(k))$ to $G_{f}^{c}(k)$, then $D(F_{f}^{c}(k))$ will form irreducible nonzero modes of $F_{f}^{c}(k)$.
@@
Then we consider a different case in which $\mathcal{C}$ ‘mixes’ a set of minimal nonzero modes $\mathcal{L}_{\varepsilon_{k}}\oplus\mathcal{L}_{-\varepsilon_{k}}$ to another set of minimal nonzero modes $\mathcal{L}_{\varepsilon^{\prime}_{k}}\oplus\mathcal{L}_{-\varepsilon^{\prime}_{k}}$. Noticing that $F_{f}^{c}(k)/F_{f}(k)\cong{Z}_{2}$, the charge conjugation $\mathcal{C}$ mixes at most two irreps of $G_{f}(k)$ to form an irreducible rep of $G_{f}^{c}(k)$, as illustrated in the Theorem 5. If the restricted rep of $D(G_{f}(k))$ from an irrep of $D(G_{f}^{c}(k))$ is reducible $D(G_{f}(k))=D_{1}(G_{f}(k))\oplus D_{2}(G_{f}(k))$, then the energy modes ±$\pm\varepsilon_{k}$ of $F^{c}_{f}(k)$ containing the restricted rep $D(G_{f}^{c}(k))$ will mix the energy modes ±$\pm\varepsilon^{\prime}_{k}$ of $F_{f}(k)$ containing the restricted rep $D_{1}(G_{f}(k))$ and the energy modes ±$\pm\varepsilon^{\prime\prime}_{k}$ containing the restricted rep $D_{2}(G_{f}(k))$. This mixing enforces the modes ±$\pm\varepsilon^{\prime}_{k}$ to be degenerate with ±$\pm\varepsilon^{\prime}_{k}$ and results in a new eigenvalues ±$\pm\varepsilon_{k}$. Namely, this will generally cause higher degeneracy. Since the structure of minimal nonzero modes remains, such $\mathcal{C}$ does yield zero modes in general. This type of charge conjugation will be encountered in SecIV.1, which causes the 2-fold degeneracy of the bands.
@@
In the above we listed the cases in which zero modes can appear due to the presence of charge conjugation symmetry $\mathcal{C}$. We call these zero modes as $\mathcal{C}$-minimal zero modes’. A natural question is, given a set of minimal nonzero modes at $k$, is it always possible to construct a $\mathcal{C}$ to turn them into $\mathcal{C}$-minimal zero modes? The answer is positive, as stated in the theorem 3. The proof of this theorem is given in the following App. G.
@@ Appendix G The proof of the theorem 3
We prove theorem 3 by illustrating the existence of $\mathcal{C}$ to turn any given minimal nonzero modes at momentum $\bm{k}$ into $\mathcal{C}$-minimal zero modes.
@@ G.1.1 When $G_{f}(k)$ is unitary
Extension: $\mathcal{C}$-minimal zero modes from reducible minimal nonzero modes $D_{+}(F_{f}(k))\oplus D_{-}(F_{f}(k))$. In this case, we have $D_{+}(\tilde{\mathcal{P}})\not\cong D_{-}(\tilde{\mathcal{P}})$ and $D_{+}(G_{f}(k))\cong D_{-}(G_{f}(k))$. According to the Frobenius reciprocity, it is not possible to do a $\mathbb{Z}_{2}$ induction such that the two reps $D_{\pm}(G_{f}(k))$ form an irreducible representation of $G_{f}^{c}(k)$. Namely, every $Z_{2}^{\mathcal{C}}$ operation preserving the given reducible minimal nonzero modes must give two extensions for the irrep $D_{\pm}(G_{f}(k))$. Moreover, in order to obtain $\mathcal{C}$-minimal zero modes, the two extensions must be non-equivalent and the $Z_{2}^{\mathcal{C}}$ operation cannot be an induction from the group $F_{f}(k)$ to $F_{f}^{c}(k)$ when restricted to the subspace of reducible minimal nonzero modes.
There are different ways to do the $Z_{2}$ extensions, we only consider the simplest one. As the effective charge-conjugation $\mathcal{C}$ commutes with all $F_{f}(k)$, we obtain the group $F_{f}^{c}(k)=F_{f}(k)\times{Z}_{2}^{\mathcal{C}}$. In the direct sum space $D_{+}(F_{f}(k))\oplus D_{-}(F_{f}(k))$, $\mathcal{C}$ is reducible. Consequently one obtains two non-equivalent extension of $D_{+}(G_{f}(k))=D_{-}(G_{f}(k))$ by the two 1D reps of ${Z}_{2}$, resulting in $D_{+}(G_{f}^{c}(k))\not\cong D_{-}(G_{f}^{c}(k))$ (here $\mathcal{C}$ provides two different quantum numbers in $D_{+}(G_{f}^{c}(k))$ and $D_{-}(G_{f}^{c}(k))$ respectively). Considering $D_{+}(\tilde{\mathcal{P}})\not\cong D_{-}(\tilde{\mathcal{P}})$, now one obtains a set of reducible minimal zero modes for the group $F^{c}_{f}(k)$. This construction is applied in several lattice models in Sec.IV.
Induction: $\mathcal{C}$-minimal zero modes from irreducible nonzero modes $D(F_{f}(k))$. In this case, the restricted rep $D(G_{f}(k))$ from the irrep $D(F_{f}(k))$ can be reduced into a direct sum of two nonequivalent irreps, namely $D(G_{f}(k))=d_{1}(G_{f}(k))\oplus d_{2}(G_{f}(k))$. It is not possible to do extension on subgroup $G_{f}(k)$ to obtain a set of minimal zero modes. But a ${Z}_{2}$ induction is straightforward if $\mathcal{C}$ is closed in the space $D(G_{f}(k))$. Then $D(\mathcal{C})$ is irreducible in the space $D(G_{f}(k))$ since $d_{1}(G_{f}(k))$ and $d_{2}(G_{f}(k))$ are non-equivalent. Thus obtained $D(G_{f}^{c}(k))$ forms a set of irreducible zero modes.
@@ G.1.2 When $G_{f}(k)$ is anti-unitary
Supposing that $D(G_{f}^{c}(k))$ form an irreducible representation of $G_{f}^{c}(k)$, then there is only one hermitian centralizer for $D(G_{f}^{c}(k))$, namely the identity matrix. If $d_{2}(G_{f}(k))\not\cong d_{1}(G_{f}(k))$, then the induced rep $D(G_{f}^{c}(k))$ is always irreducible; if $d_{2}(G_{f}(k))=d_{1}(G_{f}(k))$, to make $D(G_{f}^{c}(k))$ irreducible the following condition should be satisfied,
@@
The simplest example for an induction with $d_{2}(G_{f}(k))=d_{1}(G_{f}(k))$ is the group $G_{f}^{c}(k)=G_{f}(k)\times{Z}_{2}^{\mathcal{C}}$ with $\omega_{2}(\mathcal{C},\mathcal{C})=1$ and $L(\mathcal{C})=\mathrm{i}I$, namely $D(\mathcal{C})=\Gamma_{y}=\tau_{y}\otimes I$. Since $D(\mathcal{C})$ cannot be block diagonalized via real matrix, $D(\mathcal{C})$ and the rep matrices of antiunitary elements in $G_{f}(k)$ cannot be simultaneously block diagonalized, meaning that $D(G_{f}^{c}(k))$ is irreducible. Hence $D(F_{f}^{c}(k))$ correspond to a set of irreducible zero modes. This kind of induction is applied in the construction of lattice models in Sec.IV.
@@
$\mathcal{C}$-minimal zero modes from the reducible minimal nonzero modes $D_{+}(F_{f}(k))\oplus D_{-}(F_{f}(k))$. In this case, $D_{+}(G_{f}(k))$ and $D_{-}(G_{f}(k))$ are equivalent. One can always do inductions following the construction for the irreducible nonzero modes with the case that $d_{1}(G_{f}(k))$ is equivalent to $d_{2}(G_{f}(k))$.
Remark: We remark that, for $D_{\pm}(G_{f}(k))$ belong to the $\mathbb{C}$ class and $D_{+}(F_{f}(k))\cong D_{-}(F_{f}(k))$ [corresponding to the case (4) in the Tab.1], one can just do a ${Z}_{2}$ induction on each irreducible blocks for the subgroup $H_{f}(k)$. Then the new irreducible representation of the group $G_{f}^{c}(k)$ is of the $\mathbb{R}$ class. This also leads to a set of $\mathcal{C}$-minimal zero modes.
Moreover, one can also do extensions to obtain a set of reducible minimal zero modes following the discussions in the subsection G.1.1.

### v3

@@ F.3 Consequences of $\mathcal{C}$ symmetry
Firstly, we consider a special case in which $\mathcal{C}$ preserves the supporting space of every minimal nonzero modes ±$\mathcal{L}_{\pm\varepsilon_{k}}$, namely we first assume that $\mathcal{C}$ does not cause additional degeneracy for the energy eigenstates with $\varepsilon_{k}\not=0$. Such charge-conjugation symmetries all have the following property: for any given minimal nonzero mode, there exists an extension from $F_{f}(k)$ to $F^{c}_{f}(k)$. Recall that a set of minimal nonzero modes is either
(I) a set of irreducible nonzero modes which carries a composite irrep of $F_{f}(k)$ denoted as $D_{a}(F_{f}(k))$ with the restricted rep $D_{a}(G_{f}(k))=d_{a}(G_{f}(k))\oplus\tilde{d}_{a}(G_{f}(k))$, or
@@
(II) a set of RMNZM which form a direct sum of two simple irreps of $F_{f}(k)$ denoted as $D_{a}(F_{f}(k))=\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$).
@@
Then the presence of $\mathcal{C}$ can affect the set of minimal nonzero modes in different ways. Here we discuss them separately. From now on, we will omit the subscript a in the labels of reps $D_{a}$ and $d_{a},\tilde{d}_{a}$.
(I) For the irreducible nonzero modes $D(F_{f}(k))$.
@@
1) If the irrep $D(G_{f}^{c}(k))$ is an induction from $d(G_{f}(k))$ (or from $\tilde{d}(G_{f}(k))$), namely, there does not exist any extension from $d(G_{f}(k))$ or $\tilde{d}(G_{f}(k))$ to $G_{f}^{c}(k)$, then $D(F^{c}_{f}(k))$ corresponds a set of irreducible zero modes of $F_{f}^{c}(k)$.
@@
2) If there exists an extension from $d(G_{f}(k))$ (or $\tilde{d}(G_{f}(k)))$ to $d(G^{c}_{f}(k))$ (or $\tilde{d}(G^{c}_{f}(k))$), then $D(F^{c}_{f}(k))$ remains to be an irreducible nonzero mode of $F_{f}^{c}(k)$.
(II) For the RMNZM $D(F_{f}(k))=\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$.
@@
1) If there exist an extension from $\mathcal{D}_{\pm}(F_{f}(k))$ to $\mathcal{D}_{\pm}(F_{f}^{c}(k))$ (hence there must exist an extension from $d(G_{f}(k))$ to $d(G_{f}^{c}(k))$), then $\mathcal{D}_{\pm}(F^{c}_{f}(k))$ are two simple irreps of $F_{f}^{c}(k)$. If $\mathcal{D}_{\pm}(F^{c}_{f}(k))$ satisfy the conditions listed in Tab.1, namely if they can couple to each other, then $\mathcal{D}_{+}(F^{c}_{f}(k))\oplus\mathcal{D}_{-}(F^{c}_{f}(k))$ remains to be a set of RMNZM; on the other hand, if they cannot couple to each other, (for example, $\mathcal{C}$ provides different quantum numbers in $\mathcal{D}_{+}(F^{c}_{f}(k))$ and $\mathcal{D}_{-}(F^{c}_{f}(k))$ to prevent them from coupling to each other), then $\mathcal{D}_{+}(F^{c}_{f}(k))\oplus\mathcal{D}_{-}(F^{c}_{f}(k))$ will form reducible minimal zero modes of $F^{c}_{f}(k)$.
@@
2) If $\mathcal{C}$ couples $\mathcal{D}_{+}(F_{f}(k))$ with $\mathcal{D}_{-}(F_{f}(k))$ such that the direct sum space $\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$ forms a simple irrep $D(F_{f}^{c}(k))$ for $F_{f}^{c}(k)$ (namely $D(F_{f}^{c}(k))$ and $D(G_{f}^{c}(k))$ are both irreducible), then $D(F_{f}^{c}(k))$ forms irreducible zero modes of $F_{f}^{c}(k)$.
@@
3) If $\mathcal{C}$ couples $\mathcal{D}_{+}(F_{f}(k))$ with $\mathcal{D}_{-}(F_{f}(k))$ such that the direct sum space $\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$ forms a composite irrep $D(F_{f}^{c}(k))$ for $F_{f}^{c}(k)$ (namely $D(F_{f}^{c}(k))$ is irreducible but $D(G_{f}^{c}(k))$ is reducible), then $D(F_{f}^{c}(k))$ will form irreducible nonzero modes of $F_{f}^{c}(k)$.
@@
Secondly, we consider a different case in which $\mathcal{C}$ ‘joins’ a set of minimal nonzero modes ±$\mathcal{L}_{\pm\varepsilon^{\prime}_{k}}$ with another set of minimal nonzero modes ±$\mathcal{L}_{\pm\varepsilon^{\prime\prime}_{k}}$. Actually, since $G_{f}^{c}(k)/G_{f}(k)=F_{f}^{c}(k)/F_{f}(k)\cong{Z}_{2}$, as illustrated in theorem 5 the charge conjugation $\mathcal{C}$ joins at most two irreps of $G_{f}(k)$ to form an irreducible rep of $G_{f}^{c}(k)$, namely $\mathcal{C}$ involves at most two sets of minimal nonzero modes. The direct sum space =±$\mathcal{L}=\mathcal{L}_{\pm\varepsilon^{\prime}_{k}}\oplus\mathcal{L}_{\pm\varepsilon^{\prime\prime}_{k}}$ form a representation space of the group $F_{f}^{c}(k)$ as well as $G_{f}^{c}(k)$. If $\mathcal{L}$ contains two irreps of $G_{f}^{c}(k)$, namely if $\mathcal{L}$ carries a composite irreps of $F_{f}^{c}(k)$ or a direct sum of two simple irreps (which are couplable partner of each other) of $F_{f}^{c}(k)$, then $\mathcal{L}$ forms an enlarged set of minimal nonzero modes of $F_{f}^{c}(k)$ with doubled degeneracy; otherwise, if $\mathcal{L}$ contains four irreps of $G^{c}_{f}(k)$, then the degeneracy of the nonzero modes will not be enlarged. The $\mathcal{C}$ enlarged degeneracy occurs in Sec.IV.1, where the $\mathcal{C}$ symmetry causes the 2-fold degeneracy of the bands. Such kind of charge conjugation $\mathcal{C}$ does not yield zero modes in general.
@@
In the above discussion, we listed the cases in which zero modes can appear due to the presence of charge conjugation symmetry $\mathcal{C}$. We call these zero modes as ‘minimal $\mathcal{C}$-zero modes’, including the irreducible $\mathcal{C}$-zero modes and the reducible minimal $\mathcal{C}$-zero modes. A natural question is, given a set of minimal nonzero modes at $k$, is it always possible to construct a $\mathcal{C}$ to turn them into minimal $\mathcal{C}$-zero modes? The answer is positive, as stated in the theorem 3. The proof of this theorem is given in the following App. G.
@@ Appendix G The proof of the theorem 3
We prove theorem 3 by illustrating the existence of $\mathcal{C}$ to turn any given minimal nonzero modes of $F_{f}(k)$ at momentum $\bm{k}$ into minimal $\mathcal{C}$-zero modes. We assume that $\mathcal{C}$ preserves the space of minimal nonzero modes of $F_{f}(k)$.
@@ G.1.1 When $G_{f}(k)$ is unitary
Extension: minimal $\mathcal{C}$-zero modes from RMNZM $\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$. In this case, we have $D_{+}(\tilde{\mathcal{P}})\not\cong D_{-}(\tilde{\mathcal{P}})$ and $\mathcal{D}_{+}(G_{f}(k))\cong\mathcal{D}_{-}(G_{f}(k))$. When $\mathcal{C}$ is unitary then the induced rep from $\mathcal{D}_{+}(G_{f}(k))$ or $\mathcal{D}_{-}(G_{f}(k))$ to $G^{c}_{f}(k)$ is reducible. Namely, it is not possible to do an induction from $\mathcal{D}_{+}(G_{f}(k))$ or $\mathcal{D}_{-}(G_{f}(k))$ such that the induced irrep of $G^{c}_{f}(k)$ is equivalent to $\mathcal{D}_{+}(G_{f}(k))\oplus\mathcal{D}_{-}(G_{f}(k))$ when restricted to $G_{f}(k)$. Hence, minimal $\mathcal{C}$-zero modes can only be obtained by extensions from $\mathcal{D}_{\pm}(F_{f}(k))$ to $\mathcal{D}_{\pm}(F_{f}^{c}(k))$.
There are different ways to do the $Z_{2}$ extensions, we only consider the simplest case by assuming that the effective charge-conjugation $\mathcal{C}$ commutes with $F_{f}(k)$, namely $[\mathcal{C},F_{f}(k)]=0$. In the direct sum space $\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$, $\mathcal{C}$ is reducible. Consequently one obtains two non-equivalent extension of $\mathcal{D}_{+}(G_{f}(k))=\mathcal{D}_{-}(G_{f}(k))$ by the two 1D reps of ${Z}_{2}$, resulting in $\mathcal{D}_{+}(G_{f}^{c}(k))\not\cong\mathcal{D}_{-}(G_{f}^{c}(k))$ (here $\mathcal{C}$ provides two different quantum numbers in $\mathcal{D}_{+}(G_{f}^{c}(k))$ and $\mathcal{D}_{-}(G_{f}^{c}(k))$ respectively). Considering $\mathcal{D}_{+}(\tilde{\mathcal{P}})\not\cong\mathcal{D}_{-}(\tilde{\mathcal{P}})$, now one obtains a set of reducible minimal $\mathcal{C}$-zero modes for the group $F^{c}_{f}(k)$. This construction is applied in several lattice models in Sec.IV.
Induction: minimal $\mathcal{C}$-zero modes from irreducible nonzero modes $D(F_{f}(k))$. In this case, the restricted rep $D(G_{f}(k))$ from the irrep $D(F_{f}(k))$ can be reduced into a direct sum of two nonequivalent irreps, namely $D(G_{f}(k))=d_{1}(G_{f}(k))\oplus d_{2}(G_{f}(k))$. It is not possible to do extension on subgroup $G_{f}(k)$ to obtain a set of minimal zero modes. But a ${Z}_{2}$ induction is straightforward if $\mathcal{C}$ is closed in the space $D(G_{f}(k))$. Then $D(\mathcal{C})$ is irreducible in the space $D(G_{f}(k))$ since $d_{1}(G_{f}(k))$ and $d_{2}(G_{f}(k))$ are non-equivalent. Thus obtained $D(G_{f}^{c}(k))$ forms a set of irreducible $\mathcal{C}$-zero modes.
@@ G.1.2 When $G_{f}(k)$ is anti-unitary
Supposing that $D(G_{f}^{c}(k))$ form an irrep of $G_{f}^{c}(k)$, then there is only one hermitian centralizer for $D(G_{f}^{c}(k))$, namely the identity matrix. If $d_{2}(G_{f}(k))\not\cong d_{1}(G_{f}(k))$, then the induced rep $D(G_{f}^{c}(k))$ is always irreducible; if $d_{2}(G_{f}(k))=d_{1}(G_{f}(k))$, to make $D(G_{f}^{c}(k))$ irreducible the following condition should be satisfied,
@@
The simplest example for an induction with $d_{2}(G_{f}(k))=d_{1}(G_{f}(k))$ is the group $G_{f}^{c}(k)=G_{f}(k)+{\mathcal{C}}G_{f}(k)$ with $\mathcal{C}$ commuting with $G_{f}(k)$, $\omega_{2}(\mathcal{C},\mathcal{C})=1$ and $L(\mathcal{C})=\mathrm{i}I$, namely $D(\mathcal{C})=\Gamma_{y}=\tau_{y}\otimes I$. Since $D(\mathcal{C})$ cannot be block diagonalized via real matrix, $D(\mathcal{C})$ and the rep matrices of antiunitary elements in $G_{f}(k)$ cannot be simultaneously block diagonalized, meaning that $D(G_{f}^{c}(k))$ is irreducible. Hence $D(F_{f}^{c}(k))$ correspond to a set of irreducible $\mathcal{C}$-zero modes. This kind of induction is applied in the construction of lattice models in Sec.IV.
@@
Minimal $\mathcal{C}$-zero modes from the RMNZM $\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$. In this case, $\mathcal{D}_{+}(G_{f}(k))$ and $\mathcal{D}_{-}(G_{f}(k))$ are equivalent. One can always do inductions following the previous construction for the irreducible nonzero modes with $d_{1}(G_{f}(k))=d_{2}(G_{f}(k))$.
Remark: We comment that, for $\mathcal{D}_{\pm}(G_{f}(k))$ belong to the $\mathbb{C}$ class and $\mathcal{D}_{+}(F_{f}(k))\cong\mathcal{D}_{-}(F_{f}(k))$ [corresponding to the case (4) in the Tab.1], one can just do a ${Z}_{2}$ induction on each irreducible blocks for the subgroup $H_{f}(k)$. Then the new irrep of the group $G_{f}^{c}(k)$ is of the $\mathbb{R}$ class. This also leads to a set of reducible minimal $\mathcal{C}$-zero modes according to the case (2) in the Table.1. Moreover, one can also do extensions to obtain a set of reducible minimal $\mathcal{C}$-zero modes following the discussions in the subsection G.1.1.